Circular No. 8626

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2005 Q7–Q10, R5–R7, S5–S11 (SOHO) Additional comets have been found on SOHO website images (cf. IAUC

8616; QY = Q. Ye); all were Kreutz sungrazers except for C/2005 Q8, a Meyer-group object that appeared stellar and showed no elongation. All the objects that were visible in C3 images appeared stellar and tailless therein; peak C3 magnitudes for some of the objects: C/2005 Q7, 7.3, 12.4R_{⊙}, Aug. 25.195 UT; C/2005 Q9, 7.1, 12.2R_{⊙}, Aug. 30.388; C/2005 Q10, 7.4, 12.4R_{⊙}, Aug. 30.779; C/2005 R5, 7.4, 12.0R_{⊙}, Sept. 6.054; C/2005 R7, 6.6, 12.6R_{⊙}, Sept. 16.112; C/2005 S5, 7.7, 11.1R_{⊙}, Sept. 19.012; C/2005 S6, 6.4, 13.4R_{⊙}, Sept. 20.779; C/2005 S8, 6.4, 12.9R_{⊙}, Sept. 23.935; C/2005 S9, 7.6, 12.5R_{⊙}, Sept. 24.721; C/2005 S10, 6.8, 12.2R_{⊙}, Sept. 25.096. C/2005 R6 was too faint for photometry. C/2005 R7 was faint and diffuse with no tail in C2 images. C/2005 S5 was elongated in C2 images. C/2005 S6 was faint and diffuse with a hint of a tail in C2 images. C/2005 S7 was too faint for C3 photometry. C/2005 S5 and S7–S11 appeared very faint and diffuse in C2 images. C/2005 S11 reached mag ∼ 8 in C3 images.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 Q7 Aug. 24.571 9^{h}49.^{m}1 + 9^{o}30′ C3  SF 2005-U71
C/2005 Q8  25.246 10 11.0 +12 33 C2  HS 2005-U71
C/2005 Q9  29.696 10 06.7 + 7 55 C3  SF 2005-U71
C/2005 Q10  30.321 10 12.6 + 7 46 C3  BZ 2005-U71
C/2005 R5 Sept. 5.863 10 41.2 + 5 45 C3  HS 2005-U72
C/2005 R6 6.263 10 41.7 + 5 37 C3 HS 2005-U72
C/2005 R7  15.363 11 05.4 + 1 32 C3/2 HS 2005-U72
C/2005 S5  18.613 11 26.3 − 0 38 C3/2 HS 2005-U72
C/2005 S6  20.388 11 28.8 − 0 09 C3/2 HS 2005-U85
C/2005 S7  23.346 11 44.4 − 0 57 C3/2 QY 2005-U85
C/2005 S8  23.462 11 38.8 − 1 17 C3/2 HS 2005-U85
C/2005 S9  24.448 11 46.4 − 1 31 C3/2 HS 2005-U85
C/2005 S10  24.763 11 46.8 − 1 48 C3/2 TC 2005-U85
C/2005 S11  25.279 11 51.2 − 1 40 C3/2 HS 2005-U85

V1118 ORIONIS E. O. Waagen, AAVSO, reports that this star has remained bright this

year (cf. IAUC 8460); selected magnitude estimates, in part from Waagen: Mar. 6.451 UT, 12.7 (P. Williams, Heathcote, N.S.W.; visual); Apr. 22.388, 12.8 (Williams); Oct. 26.599, V = 14.1 (G. De Scala, Sydney, N.S.W.; CCD); 30.003, 13.7 (J.-C. Merlin, Le Creusot, France; CCD).

2005 November 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green